a bag contains 12 pieces of candy, five of which are lemon drops. If you reach in and select two pieces of candy. What is the probability they will both be lemon drops?

Answer by malakumar_kos@yahoo.com(315)


a bag contains 12 pieces of candy, five of which are lemon drops. If you reach in and select two pieces of candy. What is the probability they will both be lemon drops?
number of candy pieces in the bag = n = 12
number of candy pieces to be selected = r = 2
number of ways of selecting 2out of 12 = N(s) = 12c2 = 12*11/2 = 6*11 = 66
number of lemon drops = 5
number to be selected = 2
number of favourable events = N(A) = 5c2 = 5*4/2 = 5*2 = 10
Probability of selecting 2 lemon drops = P(A) = N(A)/N(s)
= 10/66
= 5/33